# references.bib

### v1

@@ 5 Discussion
This proof–of–concept study demonstrates that a conditional autoencoder (CAE) can recover physically meaningful latent variables from galaxy images while remaining computationally efficient. In our experiments, the CAE assigns total flux and half–light radius to two dedicated latent coordinates, $z_{1}$ and $z_{2}$, respectively. The strong one–to–one relations in Figure 3 confirm that the supervision term in Eq. (2) successfully aligns these axes with their target parameters. Manipulating each coordinate in isolation produces the expected response in image space—brightness varies smoothly with $z_{1}$, whereas the apparent size changes with $z_{2}$ (Figure 4). Such behavior indicates a disentangled latent geometry in which individual dimensions control single astrophysical factors of variation.
@@
Several caveats must be addressed before the method can be deployed in production pipelines. (i) Simulated data. Our network is trained exclusively on idealised GalSim images; real survey data contain blended sources, cosmic rays, detector artefacts and rare morphologies that can degrade performance. Domain–adaptation strategies—such as fine–tuning on a small, labeled subset of observations—will be necessary. (ii) Limited parameter set. We supervise only two parameters. Extending the framework to ellipticity, Sérsic index, bulge–to–disc ratio or point–source contamination is feasible but will require a larger latent dimensionality and stronger regularization. (iii) High–concentration profiles. The CAE occasionally under–fits highly concentrated ($n\gtrsim 3$) galaxies. Incorporating a multi–scale or parametric likelihood loss may alleviate this.

### v2

@@ 5 Discussion
This proof–of–concept study demonstrates that a conditional autoencoder (CAE) can recover physically meaningful latent variables from galaxy images while remaining computationally efficient. In our experiments, we assign total flux and half-light radius using our $z_{\mathrm{phy}}$ layer to two dedicated physical latent coordinates, $z_{1}$ and $z_{2}$, respectively. The strong one–to–one relations in Figure 3 confirm that the supervision term in Eq. (2) successfully aligns these axes with their target parameters. Manipulating each coordinate in isolation produces the expected response in image space—brightness varies smoothly with $z_{1}$, whereas the apparent size changes with $z_{2}$ (Figure 5). Such behavior indicates a disentangled latent geometry in which individual dimensions control single astrophysical factors of variation.
@@
Several caveats must be addressed before the method can be deployed in production pipelines. (i) Simulated data. Our network is trained exclusively on idealised GalSim images; real survey data contain blended sources, cosmic rays, detector artefacts and rare morphologies that can degrade performance. Hence, an important extension would be to utilise real data from surveys such as Euclid or LSST. Domain-adaptation strategies will be necessary. For example, concerning blending, a strategy than can be adopted is forcing the encoder to produce similar latent features for simulated and blended galaxies, making the decoder more robust to blending artifacts [undefaac] (ii) Limited parameter set. We supervise only two parameters. Extending the framework to ellipticity, Sérsic index, bulge–to–disc ratio or point–source contamination is feasible but will require a larger latent dimensionality and stronger regularization. (iii) High–concentration profiles. The CAE occasionally under–fits highly concentrated ($n\gtrsim 3$) galaxies. Incorporating a multi–scale or parametric likelihood loss may alleviate this.
@@
Acknowledgements
The authors would like to thank Weishun Zhong for suggesting this project and for providing guidance throughout the numerical experiments and the preparation of this manuscript.
Funding
The authors do not have any funding information to disclose.
@@ Acknowledgements
The primary objective of our autoencoder architecture is to provide architecture which works well as a proof-of-concept for future use of conditional autoencoders.